Everton College Store paid $1666 for 47 calculators. The store paid $10 for each scientific calculator. The others all graphing calculators, cost the store $56 each. How many of each type of calculator was ordered?
Answer by swincher4391(1107)

Let x be scientific, y be everything else.
x+y = 47
10x + 56y = 1666
--------------------
10x +10y = 470
10x +56y = 1666
----------------
0x -46y = -1196
y = 26
x + 26 = 47
x = 21
----------------------
21 scientific, 26 other